An experiment has five outcomes, I, II, III, IV, and V. If P (I) = 0. 08, P (II) = 0. 20, and P (III) = 0. 33, what are the possible values for the probability of outcome V? If outcomes IV and V are equally likely, what are their probability values?
Answer by richwmiller(17219)

P (I) = 0. 08
P (II) = 0. 20
P (III) = 0. 33
.08+.20+.33=0.61
1-0.61=.39

P (IV)+P (V)=.39
in second case P (IV)=P (V)
P (IV) =.195
P (V) =.195